The weight of 1000 children average 50 kg and the standard deviation is 5kg.


26. What is the z-score corresponds to 40 kg?

1
Expert's answer
2022-04-26T23:43:19-0400

"Z=\frac{x-\mu}{\sigma}=\frac{40-50}{5}=-2"